# GPU-accelerated Evolutionary Multiobjective Optimization Using Tensorized RVEA

### v3

@@ 4.1. Acceleration Performance
In this experiment, the acceleration performance of TensorRVEA is rigorously compared against the standard RVEA on the DTLZ1 problem (Deb et al., 2005). We conducted a comprehensive evaluation involving algorithms: RVEA (CPU), TensorRVEA (CPU), RVEA (GPU), and TensorRVEA (GPU), with the parenthetical notations indicating the computational devices used for each algorithm’s execution. Two sets of experiments were conducted for each algorithm, each iteration independently repeated 10 times to ensure robustness. In the first series, we standardized the decision dimension and the number of objectives at $d=100$ and $m=3$, respectively, while varying the population size from 32 ($2^{5}$) to 16,384 ($2^{14}$). The algorithms were run over 100 generations, during which we calculated the average computation time per generation. The second series maintained a constant population number ($n=100$) and the number of objectives ($m=3$), but varied the decision dimension, doubling it from 512 ($2^{9}$) to 262,144 ($2^{18}$). Again, these algorithms were run for 100 generations to ascertain the average time per generation. The primary focus of these experiments was to analyze the speedup achieved by TensorRVEA (GPU) compared to RVEA (CPU), particularly under scenarios of large populations and high-dimensional decision contexts, thus understanding acceleration performance in handling extensive computational loads and large-scale problems.
@@
As evidenced in Figure 3, a discernible trend is apparent in terms of runtime and population for both RVEA (CPU) and TensorRVEA (CPU). By contrast, a striking acceleration ratio is evident when comparing TensorRVEA (GPU) to RVEA (CPU), with the former achieving up to 1528$\times$ speedups. Moreover, the runtime for TensorRVEA (GPU) demonstrates relative stability across a wide range of population sizes, from 32 to 1,024. It is posited that beyond a population size of 1024, the hardware may encounter overhead limitations. Intriguingly, the performance curve of RVEA (GPU) is observed to intersect with that of the CPU variant. This can be attributed to RVEA’s implementation using JAX’s “fori_loop“ function, which may incur consistent overheads.
As evidenced in Figure 4, with the increasing problem dimensions, there is a corresponding rise in the time per generation for algorithms executed on the CPU. By contrast, the runtime of TensorRVEA (GPU) consistently remains lower than that of RVEA (GPU), achieving speedups of up to 1042$\times$ compared to RVEA (CPU). However, a noticeable increase in runtime is observed as the problem dimension escalates to 16,384. This surge can likely be attributed to the upper limitation of the GPU’s computing capacity reached. Upon examining the last four data points on a normal axis, it becomes evident that the slopes are very similar, indicating the consistent scalability of TensorRVEA.
@@ 4.3. Multiobjective Neuroevolution
As shown in Figure 6, TensorRVEA shows the best performance on all problems. The final average HV and average EU scores of TensorRVEA in the four environments surpass those of NSGA-II and random search. However, it is important to note that the confidence intervals for TensorRVEA tend to increase during the later phases, suggesting a potential variability in its performance outcomes. In contrast, the confidence intervals for both NSGA-II and random search remain comparatively narrow. Interestingly, the HV performance of NSGA-II in the MoHalfCheetah and MoSwimmer environments does not match the efficacy of random search. This observation points to the varying effectiveness of these algorithms across different problem settings. In the MoHopper-m3 environment, NSGA-II and RS show similar performance in the HV and the EU metrics. This resemblance could be attributed to the dominance resistance phenomenon and a swift increase in nondominated solutions. TensorRVEA consistently yields high-quality and diverse policies across all tested environments, underlining its versatility. This superior performance is attributable to its selection mechanism based on reference vectors. Unlike the non-dominated sorting method, this method facilitates evolution in the direction of the reference vectors, enabling it to effectively escape local optima. Notably, diversity is a crucial factor in robotic control tasks, catering to varied requirements such as high-speed operation at the expense of energy efficiency or the converse. In the MoSwimmer environment, NSGA-II produces more evenly distributed solutions, yet there remains room for improvement in solution quality. Meanwhile, the solutions generated by random search tend to lose diversity, with multiple solutions emerging at specific performance levels.
@@ 4.4. Extensibility
In this experiment, the extensibility of TensorRVEA was tested by employing a variety of reproduction operators. Within the framework of TensorRVEA, we integrated several representative reproduction operators, including Genetic Algorithm (GA) (Deb et al., 1996a), Particle Swarm Optimization (PSO) (Eberhart and Kennedy, 1995), Differential Evolution (DE) (Storn and Price, 1997), Competitive Swarm Optimizer (CSO) (Cheng and Jin, 2015), as well as the Random reproduction serving as the baseline for comparison. Here, CSO is picked, as it is tailored for large-scale optimization, a required feature for neuroevolution. The population size was fixed at 10,000, and each algorithm was executed over 100 generations in both the 2-objective and 3-objective MoHopper environments. The performance of these variations was evaluated in terms of HV and EU.
@@ C.3.1. Definition of Multiobjective Neuroevolution Problems
(17) $f_{2}=10*(h_{\text{curr}}-h_{\text{init}})-\sum_{i}{{a_{i}}^{2}}+C,$
@@
(19) $f_{2}=10*(h_{curr}-h_{\text{init}})+C,$
where $h_{curr}$ is the current height of the hopper and $h_{\text{init}}$ is the initial height. The third objective is control cost:
@@ C.3.2. Experimental Setup
MoHalfCheetah (0, -599.78643799)
MoHopper-m2 (0, -865.70227051)
MoSwimmer (0, -0.19898804)
MoHopper-m3 (0, 0, -1942.84301758)

### v4

@@ 4.1. Acceleration Performance
In this experiment, the acceleration performance of TensorRVEA is rigorously compared against the standard RVEA on the DTLZ1 problem (Deb et al., 2005). We conducted a comprehensive evaluation involving algorithms: RVEA (CPU), TensorRVEA (CPU), RVEA (GPU), and TensorRVEA (GPU), with the parenthetical notations indicating the computational devices used for each algorithm’s execution. Two sets of experiments were conducted for each algorithm, each iteration independently repeated 10 times to ensure robustness. In the first series, we standardized the decision dimension and the number of objectives at $d=100$ and $m=3$, respectively, while varying the population size from 32 ($2^{5}$) to 16 3841638416\,38416 384 ($2^{14}$). The algorithms were run over 100 generations, during which we calculated the average computation time per generation. The second series maintained a constant population number ($n=100$) and the number of objectives ($m=3$), but varied the decision dimension, doubling it from 512 ($2^{9}$) to 262 144262144262\,144262 144 ($2^{18}$). Again, these algorithms were run for 100 generations to ascertain the average time per generation. The primary focus of these experiments was to analyze the speedup achieved by TensorRVEA (GPU) compared to RVEA (CPU), particularly under scenarios of large populations and high-dimensional decision contexts, thus understanding acceleration performance in handling extensive computational loads and large-scale problems.
@@
As evidenced in Figure 3, a discernible trend is apparent in terms of runtime and population for both RVEA (CPU) and TensorRVEA (CPU). By contrast, a striking acceleration ratio is evident when comparing TensorRVEA (GPU) to RVEA (CPU), with the former achieving up to 1528$\times$ speedups. Moreover, the runtime for TensorRVEA (GPU) demonstrates relative stability across a wide range of population sizes, from 32 to $1024$.
It is posited that beyond a population size of $1024$, the hardware may encounter overhead limitations. Notably, the performance curve of RVEA (GPU) is observed to intersect with that of the CPU variant. This can be attributed to RVEA’s implementation using JAX’s “fori_loop“ function, which may incur consistent overheads.
As evidenced in Figure 4, with the increasing problem dimensions, there is a corresponding rise in the time per generation for algorithms executed on the CPU. By contrast, the runtime of TensorRVEA (GPU) consistently remains lower than that of RVEA (GPU), achieving speedups of up to 1042$\times$ compared to RVEA (CPU). However, a noticeable increase in runtime is observed as the problem dimension escalates to 16 3841638416\,38416 384. This surge can likely be attributed to the upper limitation of the GPU’s computing capacity reached. Upon examining the last four data points on a normal axis, it becomes evident that the slopes are very similar, indicating the consistent scalability of TensorRVEA.
@@ 4.3. Multiobjective Neuroevolution
As shown in Figure 6, TensorRVEA shows the best performance on all problems. The final average HV and average EU scores of TensorRVEA in the four environments surpass those of NSGA-II and random search. However, it is important to note that the confidence intervals for TensorRVEA tend to increase during the later phases, suggesting a potential variability in its performance outcomes. In contrast, the confidence intervals for both NSGA-II and random search remain comparatively narrow. The performance of NSGA-II, as measured by HV in the MoHalfCheetah and MoSwimmer environments, is inferior to that achieved by random search. This observation points to the varying effectiveness of these algorithms across different problem settings. In the MoHopper-m3 environment, NSGA-II and RS show similar performance in the HV and the EU metrics. This resemblance could be attributed to the dominance resistance phenomenon and a swift increase in nondominated solutions. TensorRVEA consistently yields high-quality and diverse policies across all tested environments, underlining its versatility. This superior performance is attributable to its selection mechanism based on reference vectors. Unlike the non-dominated sorting method, this method facilitates evolution in the direction of the reference vectors, enabling it to effectively escape local optima. Notably, diversity is a crucial factor in robotic control tasks, catering to varied requirements such as high-speed operation at the expense of energy efficiency or the converse. In the MoSwimmer environment, NSGA-II produces more evenly distributed solutions, yet there remains room for improvement in solution quality. Meanwhile, the solutions generated by random search tend to lose diversity, with multiple solutions emerging at specific performance levels.
@@ 4.4. Extensibility
In this experiment, the extensibility of TensorRVEA was tested by employing a variety of reproduction operators. Within the framework of TensorRVEA, we integrated several representative reproduction operators, including Genetic Algorithm (GA) (Deb et al., 1996a), Particle Swarm Optimization (PSO) (Eberhart and Kennedy, 1995), Differential Evolution (DE) (Storn and Price, 1997), Competitive Swarm Optimizer (CSO) (Cheng and Jin, 2015), as well as the Random reproduction serving as the baseline for comparison. Here, CSO is picked, as it is tailored for large-scale optimization, a required feature for neuroevolution. The population size was fixed at 10 0001000010\,00010 000, and each algorithm was executed over 100 generations in both the 2-objective and 3-objective MoHopper environments. The performance of these variations was evaluated in terms of HV and EU.
@@ C.3.1. Definition of Multiobjective Neuroevolution Problems
(17) $f_{2}=10\cdot(h_{\text{curr}}-h_{\text{init}})-\sum_{i}{{a_{i}}^{2}}+C,$
@@
(19) $f_{2}=10\cdot(h_{\text{curr}}-h_{\text{init}})+C,$
where $h_{\text{curr}}$ is the current height of the hopper and $h_{\text{init}}$ is the initial height. The third objective is control cost:
@@ C.3.2. Experimental Setup
MoHalfCheetah (0, --599.78643799)
MoHopper-m2 (0, --865.70227051)
MoSwimmer (0, --0.19898804)
MoHopper-m3 (0, 0, --1942.84301758)